OpenStudy (anonymous):

Need help solving for r! 6^(-r-1)=216

OpenStudy (anonymous):

would this be 6^(-r-1)=6^3?

OpenStudy (anonymous):

and then -r-1=3?

OpenStudy (mathmale):

Note that you have 6 as base on both sides of your initial equation. If the bases are the same, as they are here, then you may equate the exponents, obtaining the much simpler equation -r-1=3, exactly as @ali1029 has already stated. Solve this for r. In your shoes, I'd add r to both sides, and then subtract 3 from both sides. Go for it.

OpenStudy (anonymous):

i got r= -4 thank you very much for explaining!! :)